Drennan Corp. is growing quickly. Dividends are expected to grow at a rate of 25 percent for the next three years, with the growth rate falling off to a constant 6 percent thereafter.

Required:

If the required return is 13 percent and the company just paid a $2.50 dividend, what is the current share price? (Hint: Calculate the first four dividends.) (Do not round intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16).)

  Current share price $   
Answer #1

D1=(2.5*1.25)=3.125

D2=(3.125*1.25)=3.90625

D3=(3.90625*1.25)=4.8828125

Value after year 3=(D3*Growth Rate)/(Required rate-Growth Rate)

(4.8828125*1.06)/(0.13-0.06)

=$73.93973214

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=3.125/1.13+3.90625/1.13^2+4.8828125/1.13^3+$73.93973214/1.13^3

=$60.45(Approx).